A homeowner is covering his patio with carpet squares, the area of the patio is 100 feet square, if each carpet square measures 8 inches by 8 inches, how many squares will it take to cover the patio?
Answer by mananth(16946)

Area of patio = 100 sq.ft
square carper is 8 by 8 inches
..
since area of patio is in sq. ft. convert the length of carpet to feet.
8" = 8/12 feet
=3/4 ft.
area of carpet = 3/4 * 3/4 = 9/16 sq.ft.
number of carpet squares = area of aptio/area of carpet squares
100 / (9/16)
100*16/9
=1600/9
=225 squares